You are investigating two traits in fruit flies controlled by one autosomal dominant gene each. The B gene controls for body color (B=black or b=brown), while the r gene controls for body sculpture (R=rugose or r=smooth). After crossing heterozygous rugose black flies with homozygous smooth brown flies, you obtain an F1 generation of 100% rugose black flies.

1. What are the dominant alleles?

2. What are the genotypes of the two parent flies to perform a dihybrid cross?

3. Of the two parent flies, identify the genotype of the tester fly?

Phenotype: Observed: Expected:

black, rugose 42

black, smooth    504

brown, rugose   327

brown, smooth 82

total     1000 1000

Fill in the column labeled “Expected” with the expected number of offspring for each listed phenotype.

  1. What phenotype represents the recombinant offspring?  

Calculate the relative genetic distance between the body color gene, and the body sculpting gene—show your work, including any applicable units!  
Answer #1

Answer 1) B and R allele are dominant allele which encode Black and rugose respectively. Because the F1 progeny have Black and rugose. This suggests that dominant genes are expressed here.

Answer2)

Genotype of parents:

Parent 1 is heterozygous for two genes. That's why genotype is BbRr

Parent 2 is homozygous recessive for two genes bbrr

Answer 3)

The tester parent is that parent which is used to test the genotype of the parent in test cross.

That's why the genotype of tester parent is BbRr.

Answer 4)

Recombiant offsprings are black rugose and brown smooth. Because the lower number indicates the recombinant type.

Answer 5)

Genetic distance is calculated by

(Recombinant/total) ×100= (124/1000) ×100=12.4 centiMorgans.